a candy store wishes to mix two types of candy to get 5 pounds of $2 per pound candy. if candy A costs $3.20 per pound and candy B costs $1.70 per pound, how much of each type should they use?
Answer by robertb(5830)

3.20A + 1.70B = 2*5 = 10
A + B = 5 ==> B = 5 - A ==> 3.2A + 1.70*(5 - A) = 10
==> 3.2A + 8.5 - 1.70A = 10
==> 1.50A = 1.5
==> A = 1 pound for A
==> B = 4 pounds for B.